OpenStudy (anonymous):

solve the equation (1/3)^(3x+6) = 9x-5

OpenStudy (anonymous):

Is that really the equation?

OpenStudy (anonymous):

yeeep...\[(\frac{ 1 }{ 3 } )^{3x+6} = 9x-5\]

hartnn (hartnn):

then it does not have any closed form...can't be solved http://www.wolframalpha.com/input/?i=%281%2F3%29%5E%283x%2B6%29+%3D+9x-5

OpenStudy (anonymous):

apparently the answer is x=4/5

OpenStudy (anonymous):

i just dont know how to get there

hartnn (hartnn):

with that equation, not possible. make sure u have typed the question correctly, i would suggest to take the screenshot.

OpenStudy (anonymous):

it's from a textbook..

OpenStudy (anonymous):

@refusetofail, ur equation is correct it's an exponential equation

OpenStudy (calculusfunctions):

@refusetofail @hartnn was right. You did not write the question clearly. Since now you're offline I will just do this one. \[\left( \frac{ 1 }{ 3 } \right)^{3x +6}=9^{x -5}\]\[(3^{-1})^{(3x +6)}=(3^{2})^{(x -5)}\]\[3^{-3x -6}=3^{2x -10}\]The bases are equal, thus equate the exponents and solve for x. Thus\[-3x -6=2x -10\]Adding 10 and 3x to both sides yields\[10-6=2x +3x\]\[4=5x\]Dividing both sides by 5 yields\[x =\frac{ 4 }{ 5 }\] And that is the answer thou protested!

hartnn (hartnn):

still have doubt with that ^ ?

OpenStudy (anonymous):

omg, thankfully I figured this one out on my own at about 3am this morning, because the exact same one was on my final!